OpenStudy (thalia):

how do i solve -2(3x+2)=2x+6

OpenStudy (anonymous):

-6x - 4 = 2x + 6 -6x - 2x -4 -6 = 0 -8x -10 = 0 -8x = 10 ----------- -8 x= 10/-8